a/40=30/100
i tried and i got 12. i dont know if its the right answer can you explain it to me?

Answer by stanbon(75887)

a/40=30/100
----
Multiply both sides by 40 to get:
a = 40(30/100)
---
a = 40(3/10)
a = 120/10
a = 12
---
You are correct.
Cheers,
Stan H.